# PointOBB-v3: Expanding Performance Boundaries of Single Point-Supervised Oriented Object Detection

### v1

@@ 5.4 Ablation Studies
TABLE VIII: Ablation studies of the grouping type used in the SSC loss, and the generation range of point annotation on DIOR-R. mAP is reported by training the Oriented R-CNN.
@@
5.4.1 Two-Stage Framework
The effect of SSC, SSFF, and DS. Tab. IV examines the impact of the three proposed key components: the SSC loss, the SSFF module, and the DS matching strategy. As illustrated in Tab. IV, each of these elements plays a role in enhancing performance. This demonstrates that the scale-consistency-based SSC loss and the scale-aware SSFF module have a substantial impact, while the DS matching strategy further improves performance by addressing misalignment caused by unknown scales.
The effect of different MIL losses. Tab. VI highlights the impact of applying MIL loss across various views. The results indicate that both enhanced views contribute to accuracy improvements, likely due to their representation of distinct types of data augmentation.
The effect of the angle branch’s setting. We investigate the effect of gradient backpropagation on the angle branch within the angle acquisition module. As presented in Tab. VII, optimizing the angle alongside the base network (”Attachment”) results in quicker convergence and improved accuracy, highlighting the importance of collaborative optimization between angle and scale.
The effect of the burn-in steps. Building on the “Attachment” setting, we examine the influence of burn-in steps through MIL loss. In the “Plainly” strategy, burn-in steps are set to 0, meaning angle learning is introduced from the start. Conversely, the “Gradually” strategy applies the defined burn-in steps progressively. As indicated in Tab. VII and Tab. IX, the “Plainly” strategy disrupts the initial optimization of the MIL network, leading to reduced performance. In contrast, the “Gradually” strategy incorporates two burn-in steps after scale and angle learning have mostly converged, enabling the network to optimize progressively and achieve better results.
TABLE IX: Ablation studies of the burn-in steps. The terms step1 and step2 refer to burn-in step1 and burn-in step2 in the progressive multi-view switching strategy, respectively. The results are evaluated on the DIOR-R dataset. mAP is reported by training the Oriented R-CNN.
@@ 5.4.1 Two-Stage Framework
53.53 38.33
55.59 39.59
57.68 41.82
TABLE X: The effects of SSFF, and IAW in DIOR-R and DOTA-v1.0 datasets. We evaluate the end-to-end detector on the DIOR-R testing set and the DOTA-v1.0 testing set.
@@
The effect of point range. The right section of Tab. VIII demonstrates the effect of varying ranges during point label generation. Adding appropriate noise shows benefits over using only the center point label (i.e., 0%). This is likely because, for certain categories, relying solely on the center point limits the network’s perceptual range, reducing its ability to recognize object boundaries effectively.
The effect of grouping type in SSC loss. The left section of Tab. VIII evaluates the influence of grouping scores by scale, ratio, and proposal within the SSC loss. The results confirm the effectiveness of scale-based grouping, supporting the initial objective of our design.
@@ 5.4.2 End-to-End Framework
The effect of SSFF and IAW. As shown in Tab. X, the proposed SSFF and IAW also play a significant role in the end-to-end version of PointOBB-v3. Compared to not using them, our model’s performance improves by 11.84% on the DIOR dataset and 16.38% on the DOTA-v1.0 dataset. This demonstrates that SSFF and IAW are indispensable components of our proposed method.
@@ 6 Conclusions
Future work. There still remains a gap between PointOBB-v3 and the RBox-supervised OOD, particularly for specific categories (e.g. BR and OP). This can be attributed to their relatively unclear boundary definitions and extreme aspect ratios. This highlights the need for deeper investigation and better use of object features in aerial imagery, particularly spatial and contextual information.

### v2

@@ 5.2.1 Results on DIOR-R
5.3 Results on More Datasets
The results are displayed in Tab. 4. On the more challenging DOTA-v1.5 dataset, PointOBB-v3 exhibits a comparable trend, outperforming PointOBB by 23.13%/27.16% and surpassing PointOBB-v2 by 3.2%/6.07% in the pseudo-generation track. Similarly, on DOTA-v2.0, PointOBB-v3 achieves a 17.99%/18.57% improvement over PointOBB and a 2.88%/1.46% gain over PointOBB-v2. Our method also demonstrates strong performance on various datasets such as FAIR1M, STAR and RSAR. Specifically, PointOBB-v3 achieves improvements of 7.16%/3.66%/8.80% compared to PointOBB and 4.99%/3.85%/3.61% over PointOBB-v2 with the Rotated FCOS detector.
On the other hand, the end-to-end version of PointOBB-v3 also delivers strong performance on these more challenging datasets. Specifically, it outperforms Point2RBox-RC on DOTA-v1.5, DOTA-v2.0, FAIR1M, and STAR datasets. Even when compared to Point2RBox-SK, which incorporates human prior knowledge, our method also demonstrates highly competitive performance, particularly achieving a certain degree of superiority on the DIOR, DOTA-v1.0, and DOTA-v1.5 datasets (37.60% vs. 27.34%, 41.29% vs. 40.27%, 31.25% vs. 30.51%). Moreover, on the STAR dataset, our end-to-end version outperforms PointOBB (FCOSR) and PointOBB-v2 (FCOSR) by 2.12% and 2.31%, respectively, while achieving gains of 2.04% and 4.28% on the RSAR dataset. It is evident that the end-to-end version still has a certain performance gap compared to the two-stage version. The potential reasons for this gap are as follows: (1) In the end-to-end version, the detection branch and the MIL branch are jointly trained. However, the training objectives of these two branches differ, which can lead to conflicting gradients during optimization, ultimately affecting overall performance. (2) Additionally, in the end-to-end version, the supervision signal for the detection branch is less stable compared to the two-stage version. This instability in the supervision signal can cause less consistent convergence during training, making it more challenging for the network to reach optimal performance.
5.4 Ablation Studies
5.4.1 Two-Stage Framework
The effect of SSC, SSFF, and DS. Tab. 5 examines the impact of the three proposed key components: the SSC loss, the SSFF module, and the DS matching strategy. As illustrated in Tab. 5, each of these elements plays a role in enhancing performance. This demonstrates that the scale-consistency-based SSC loss and the scale-aware SSFF module have a substantial impact, while the DS matching strategy further improves performance by addressing misalignment caused by unknown scales.
The effect of different MIL losses. Tab. 7 highlights the impact of applying MIL loss across various views. The results indicate that both enhanced views contribute to accuracy improvements, likely due to their representation of distinct types of data augmentation.
The effect of the angle branch’s setting. We investigate the effect of gradient backpropagation on the angle branch within the angle acquisition module. As presented in Fig. 6(a) and Tab. 7, optimizing the angle alongside the base network (“Attachment”) results in quicker convergence and improved accuracy compared to the strategy without “Attachment”, highlighting the importance of collaborative optimization between angle and scale.
The effect of the burn-in steps. Building on the “Attachment” setting, we examine the influence of burn-in steps through MIL loss. In the “Plainly” strategy, burn-in steps are set to 0, meaning angle learning is introduced from the start. Conversely, the “Gradually” strategy applies the defined burn-in steps progressively. As indicated in Fig. 6(b) and Tab. 7, the “Plainly” strategy affects the initial optimization of the MIL network, resulting in performance degradation. In contrast, the “Gradually” strategy incorporates two burn-in steps after scale and angle learning have mostly converged, enabling the network to optimize progressively and achieve better results. Furthermore, as shown in Tab. 10, we experimented with multiple burn-in step settings and observed that the model’s performance is significantly affected by these variations. Notably, the settings of “6” and “8” proved to be more effective in facilitating the optimization of the model. And the settings of “4”, “8” and “8”, “10” may have led to suboptimal performance of the network due to insufficient scale or angle learning.
The effect of point range. The right section of Tab. 8 demonstrates the effect of varying ranges during point label generation. Adding appropriate noise shows benefits over using only the center point label (i.e., 0%). This is likely because, for certain categories, relying solely on the center point limits the network’s perceptual range, reducing its ability to recognize object boundaries effectively.
The effect of grouping type in SSC loss. The left section of Tab. 8 evaluates the influence of grouping scores by scale, ratio, and proposal within the SSC loss. The results confirm the effectiveness of scale-based grouping, supporting the initial objective of our design.
Table 8: Ablation studies of the grouping type used in the SSC loss, and the generation range of point annotation on DIOR-R. mAP is reported by training the Oriented R-CNN.
@@ 5.4.1 Two-Stage Framework
Table 9: Ablation studies of the burn-in steps. The terms step1 and step2 refer to burn-in step1 and burn-in step2 in the progressive multi-view switching strategy, respectively. The results are evaluated on the DIOR-R dataset. mAP is reported by training the Oriented R-CNN.
@@
0 0 53.53 38.33
0 8 55.59 39.59
6 8 57.68 41.82
4 8 54.69 39.42
8 10 53.91 38.22
Table 10: The effects of SSFF, and IAW in DIOR-R and DOTA-v1.0 datasets. We evaluate the end-to-end detector on the DIOR-R testing set and the DOTA-v1.0 testing set.
@@ 5.4.2 End-to-End Framework
The effect of SSFF and IAW. As shown in Tab. 10, the proposed SSFF and IAW also play a significant role in the end-to-end version of PointOBB-v3. Compared to not using them, our model’s performance improves by 11.84% on the DIOR dataset and 16.38% on the DOTA-v1.0 dataset. This demonstrates that SSFF and IAW are indispensable components of our proposed method.
Table 11: Comparison of computational efficiency on the DIOR-R dataset, including training epochs, training hours, number of parameters, and inference speed. “FPS” indicates the inference speed of the trained model in generating pseudo-labels. For the two-stage methods, the values in “()” and the mAP represent the detector’s performance (FCOSR).
Method Epochs Train hours Params FPS mAP50
[…]
Point2RBox-SK 12 3.78 52.92M 33.6 27.34
Point2RBox-SK 24 7.56 52.92M 33.6 28.50
Point2RBox-SK 36 11.34 52.92M 33.6 28.90
PointOBB-v3 (e2e) 24 13.55 46.58M 33.6 37.60
[…]
PointOBB 24 (12) 11.70 (2.48) 41.80M (31.93M) 26.1 (33.6) 37.31
PointOBB-v2 12 (12) 2.67 (2.48) 32.59M (31.93M) 5.9 (33.6) 34.35
PointOBB-v2 24 (12) 5.34 (2.48) 32.59M (31.93M) 5.9 (33.6) 25.88
PointOBB-v3 (two stage) 24 (12) 14.75 (2.48) 41.80M (31.93M) 23.7 (33.6) 40.18
Figure 7: Qualitative analysis on struggling cases. On the left side of each image is the ground truth, and on the right side is the prediction result.
@@ 6 Conclusions
Future work. There still remains a gap between PointOBB-v3 and RBox-supervised OOD, particularly for specific categories such as BR and OP, which are challenged by extreme aspect ratios. (as shown in Fig. 7) The long, narrow shapes of these objects make them highly sensitive to small changes in angle. In addition, categories like GF and TS struggle due to unclear or ambiguous object boundaries, making accurate detection difficult. This highlights the need for further investigation into the use of spatial and contextual features in aerial imagery to address these challenges.
Data Availability
The datasets used in this paper are publicly available. Their names and links are as follows:
Acknowledgments
This work was supported by the National Natural Science Foundation of China under Grants (42371321; 42030102) and the Ant Group.